# Mergers of compact objects with cores of massive stars: evolution, r-process and multi-messenger signatures

### v1

@@ ACKNOWLEDGMENTS
I thank Noam Soker for invaluable discussions on the physics of these mergers and their respective outcomes over the years. I thank Brian Metzger and Hagai Perets for eye-opening conversations about the accretion process during CEE. I thank Azalee Bostroem and Manisha Shrestha for helpful input on the observational properties of these systems. I thank Mathieu Renzo for insightful suggestions and discussions that improved this manuscript. I acknowledges support from the Steward Observatory Fellowship in Theoretical and Computational Astrophysics, the IAU-Gruber Fellowship, and the CHE Fellowship.

### v2

@@ ACKNOWLEDGMENTS
I thank Noam Soker for invaluable discussions on the physics of these mergers and their respective outcomes over the years. I thank Brian Metzger and Hagai Perets for eye-opening conversations about the accretion process during CEE. I thank Azalee Bostroem and Manisha Shrestha for helpful input on the observational properties of these systems. I thank Mathieu Renzo for insightful suggestions and discussions that improved this manuscript, as well as the anonymous referee for their valuable comments and recommendations that further improved its quality. I acknowledges support from the Steward Observatory Fellowship in Theoretical and Computational Astrophysics, the IAU-Gruber Fellowship, and the CHE Fellowship.